# Introspective Growth: Automatically Advancing LLM Expertise in Technology Judgment

### v2

@@ 5 Results
We evaluate whether self-questioning with automatic answer retrieval improves LLMs’ understanding of technology (5.1). We then examine why this strategy works and how it helps us understand the way internal and external knowledge are organized in LLMs – does it aid in activating latent knowledge, offset compressed representations, or primarily rely on external knowledge (5.2). Next, we compare it to CoT prompting to assess whether LLMs can independently construct sufficient background knowledge for high task performance (5.3). We analyze how model size affects question quality and judgment accuracy (5.4).
5.1 Scientific self-questioning works well for boosting LLM understanding
We sample 1000 random pairs and use them as the testbed for demonstration. We gradually add scientific QA pairs from 1-6 for each patent in a pair to the original pair of abstracts, and then have the LLM re-judge. As shown in Figure 4, the baseline represents judge accuracy, confidence, and consistency without any self-questioning or scientific retrieval. The placebo condition uses the same question but adds a randomly selected piece of text from the same paper as the found answers, matched in length to the actual answers. We observe that as more QAs from science are added, judge accuracy significantly increases and becomes better than placebo answers. The confidence of both correct and incorrect answers decreases, but the drop is significantly greater for incorrect answers. Judgments of correct answers also become noticeably more consistent, whereas judgments of incorrect answers become less consistent. Interestingly, even the placebo answer — though merely a random, same-length excerpt from the paper — enhances understanding to some extent. However, unlike actual scientific answers, its effect does not scale with the number of questions meaning there is some initial signal that helps judgment but LLM does not learn more over more rounds of relatively unrelated science. This is likely because: (1) the question alone already facilitates understanding (we will explore this further in the next experiment 5.2), and (2) the placebo text still contains scientific content, which can contribute to comprehension.
Figure 4: More science-QA improves the judge over the placebo answer: For panel (a): $p_{q_{1}}$=0.0868, $p_{q_{2}}$=0.9270, $p_{q_{3}}$=0.0934, $p_{q_{4}}$=0.0000, $p_{q_{5}}$=0.0000, $p_{q_{6}}$=0.0021.
@@ 5.2 LLMs do not know what they know
Figure 5: LLM does not know what it has known: Self-talk helps improve judgment, but not as much as scientific QA. $p$-values: $p_{q_{1}}$ to $p_{q_{5}}$ = 0.000, $p_{q_{6}}$ = 0.013 (self vs. scientific answers).
Why does attaching question–answer pairs help? If the model already possesses the relevant knowledge and reliably accesses it when making a judgment, then self-generated questions and answers should have little effect. To explore this, we test whether performance gains stem from organizing internal knowledge, recovering underutilized representations, or introducing new information (Figure 5). We compare three conditions. In the first, the model sees only the questions it generated for each patent, without answers. In the second, it sees both its self-generated questions and answers, without any externally retrieved knowledge (i.e., what it already “knows”, we term this as "self-talk"). In the third, it sees the same questions along with scientific answers retrieved from external sources (the scientific QAs).
Comparing condition one to the raw baseline (blue line), we find that even presenting the model with only its own questions improves judgment, suggesting that structured prompting activates latent knowledge and that the LLM can utilize these textual cues like a “syllabus,” helping it parse complex texts. This effect, however, does not grow with more questions, indicating that the activation is limited. Condition two probes further and examines whether the model already contains the answers in its internal database. The performance improves further by feeding self-talk, and scales with the number of questions, indicating that the model has useful internal knowledge for the task but underutilizes it unless explicitly prompted to retrieve it (orange line). This suggests that the model does not know what it knows in terms of effectively probing internal knowledge  [^4] Another version of “knowing” is more clear-cut – evaluating the validity of an LLMs own claims (true or false) and predicting which questions they will be able to answer correctly [27]. and needs guidance to fully utilize its internal information sources. Finally, condition three (green line) yields the strongest improvement, showing that while the model benefits from questioning and self-generated answers, externally grounded answers still provide a more reliable basis for understanding. The model has often seen the answer content during training, and does not necessarily need to derive it as completely new knowledge through scientific retrieval [56]. In principle, the model already “knows” these facts. (See Appendix F for further experiments of external science acquired through scientific retrieval versus compressed prior known information: We do not find meaningful differences between embedding areas of “content” unleashed through self-talk and retrieved scientific answers. Nevertheless, they are different in other meaningful ways as demonstrated through permutation tests). In summary, model internal representations are highly compressed and do not retain the full richness of information present in the original texts on which they were trained, especially if it is sparsely distributed across the corpus, as with science [52].
5.3 Self-talk and chain-of-thought converge
In our experiments, we explored the “self-talk” QA setup (without scientific retrieval), and we show it works for improving understanding (see Figure 5). This setup can be viewed as a variant of the CoT prompting strategy. To better understand the connection between self-talk and CoT prompting, we conducted two CoT-based experiments: Zero-shot CoT, where we append the instructional prompt “Think step by step, from components to concepts” following the method of [62, 45, 28], and Few-shot CoT, where we provide the model with two annotated examples — one showing a pair of patents that are rephrasing and one showing a pair that is superficially similar but different. We compare performance across four conditions (Figure 6): a single self-talk QA, six rounds of self-talk QA, zero-shot CoT, and few-shot CoT. Our findings reveal that for small-to-mid-sized models, such as LLaMA 8B, zero-shot CoT fails to meaningfully activate reasoning capabilities. While few-shot CoT shows modest improvements, its gains are significantly outpaced by the multi-step self-talk QA approach, as shown in Figure 6 (a). To explore this further we prompted the Llama-8B model to output its few-shot CoT reasoning text across steps explicitly. The model produced $6.70\pm 0.11$ CoT reasoning steps on average. We performed a cumulative-addition experiment where we preserved the original order of the first six CoT steps, and progressively appended QAs one at a time (from 1 to 6) in a random order. At each step $k$, we computed the text similarity between the first $k$ CoT steps and the first $k$ randomized QA texts. As shown in Figure 6 (b), the CoT reasoning and self-talk QAs become semantically aligned and converge across reasoning steps. The initial steps of CoT reasoning are likely more general, but as steps accumulate, CoT becomes increasingly specific and focused, gradually forming stronger semantic connections with the self-talk QA content.
Figure 6: Few-shot CoT and self-talk QAs manifest different performances with self-talk QAs performing better (panel a), but both boost understanding and eventually converge (panel b).
5.4 Questions generated by smaller models are compatible with larger models
@@
As larger models show a higher ability to differentiate, does that also imply they generate better questions? The answer depends on which model is doing the answering. We test two models—LLaMA 3B and 8B—on questions generated by four others: LLaMA 1B, 8B, 70B, and 405B. We select 3B and 8B as answerer models because the 1B model produces low-confidence responses (94.4% with a confidence score of 0), indicating little understanding of content. In contrast, the 70B and 405B models already achieve strong performance without external scientific context, leaving little room for improvement. The 3B and 8B models provide a more informative middle ground: they show some understanding but still benefit from superior prompts.
@@
As shown in Figure 7 panels a and b, we find that the 3B model performs best when answering questions from the 1B model, and the 8B model benefits from questions posed by both the 1B and 8B. This asymmetry reveals a critical insight: questions from larger models do not consistently help smaller models, and often introduce confusion. A more knowledgeable model cannot fully substitute for a smaller model’s understanding, suggesting that knowledge extraction is not smoothly transferable across model scales [33]. At the same time, we observe that questions from smaller models often align well with the answering capacity of larger models. This may be because larger models can readily interpret, build on, and give a better answer to smaller models’ questions. Large model questions appear to go beyond the bounds (and dimensionality) of smaller model understanding and answering capacity.
@@ 6 Final Remarks
Limitations: see Appendix H. Conclusions: Self-questioning not only enhances LLMs’ understanding but also reveals how they organize and access internal/external knowledge. While models hold useful internal insights, these are often latent and underused—even CoT prompting cannot fully unlock them. Self-questioning helps activate this hidden compressed capacity, but falls short for complex tasks. Interestingly, the fundamental questions from smaller models often align with the reasoning needs of larger models. This suggests a scalable strategy: use smaller models to structure problems and guide deeper reasoning in stronger models. Impact: our work and the associated new benchmark on patent distinction offer broad insights into how AI can be used to automate the evaluation of scientific and technical contributions.
Acknowledgements
We thank Dr. Ari Holtzman, Sida Li, Jingyang Yi, and Jiazheng Wang from UChicago Data Science Institute and members from UChicago Knowledge Lab for their feedback. This work is completed with the resources provided by the UChicago Research Computing Center. We conducted all experiments using two NVIDIA A100 GPUs (40 GB each) and 150 GB of system RAM.
@@ Appendix A Understanding Through Differentiation
Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures; [18] it fundamentally involves grasping the relationships that connect concepts and principles within a domain. This requires both an appreciation of and the ability to uncover underlying structures, such as causal links, correlations, hierarchies, or analogical parallels, which distinguish true comprehension from sophisticated pattern matching or associative recall. Evaluating an LLM’s understanding, therefore, requires methods that probe its capacity to navigate these relational structures, moving beyond tests of surface-level knowledge.
@@
Cognitive science, from [16] through [54] and Gentner’s structure-mapping work [35] to recent relational-reasoning research [1], demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”. Differentiation is the capacity to distinguish concepts, entities, or situations, recognizing the features or relations that make them distinct – essentially, identifying what makes ’A’ different from ’B’. Without this ability to draw meaningful distinctions, even simple relationships cannot be properly conceptualized. Generalization, conversely, involves identifying commonalities across different instances, abstracting shared properties or underlying structures, and treating distinct items as equivalent for certain purposes, such as belonging to the same category or representing the same concept. These processes are fundamental not just for forming basic concepts, but for constructing the complex relational knowledge that constitutes deep understanding in any domain.
@@
Moreover, the ability to differentiate and generalize remains vital for constructing and refining our understanding of relationships between concepts, especially causal ones, which serve as the backbone of understanding. To explain why something occurs, we must first detect patterns: when does outcome $B$ follow event $A$, and when does it not? This requires the capacity to distinguish $A$ from alternatives like $A^{\prime}$, and to notice that $B$ follows $A$ but not $A^{\prime}$. Such contrasts form the basis for counterfactual reasoning, allowing us to imagine what would happen if everything about $A$ and $A^{\prime}$ were held constant except for one defining feature—if changing that feature flips the outcome, it signals a likely cause. At the same time, we must also recognize when observations are similar in respects that matter. Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes $B$.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims. Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation [42, 30, 22]. Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes [17].
@@
LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language [57], [58], [38], [8]. Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model [55] that describes and explains how things relate to one another. The discussion above highlights that for LLMs to demonstrate comprehension of high-level concepts, they must be able to discern similarities and differences between complex concepts. Consequently, measuring an LLM’s proficiency in discerning similarity and difference within highly specific concepts and fields offers a powerful proxy for evaluating its understanding of the conceptual foundations and landscape of these fields.
@@
Our measure of understanding draws inspiration from previous benchmarks used to assess language comprehension in earlier generations of language models. GLUE [57], for example, includes tasks that involve determining whether two sentences convey similar meanings or whether one entails the other. Similarly, our tasks focus on whether models can distinguish subtle conceptual differences. While prior benchmarks often relied on synthetic or crowd-sourced data, e.g [7], our use of patent abstracts introduces a more challenging and high-stakes setting, requiring fine-grained reasoning over technical content.
@@ Appendix B Related Literature
Relations to other benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding. But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after [36, 64, 12]. High performance may reflect superficial pattern matching or memorization rather than true conceptual comprehension. In particular, current evaluation datasets frequently allow models to exploit statistical associations or training data overlaps to get the right answers without genuine understanding  [^5] https://www.sciencenews.org/article/ai-understanding-reasoning-skill-assess. Even when test content is novel, models can exploit formatting quirks. Alzahrani et al. [2] showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions. Such brittleness indicates models are learning task-specific tricks (e.g., defaulting to a particular choice position or prompt wording) rather than demonstrating robust knowledge application. Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics [3]. Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities [9]. In sum, many current benchmarks are static and shortcut-prone, failing to truly assess whether an AI understands concepts or is simply leveraging surface regularities. Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase [4] or contrast and counterfactual sets [31]  [^6] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content [63]..
@@
Organizing internal and external knowledge: A core challenge in LLM reasoning is how to activate the vast latent knowledge embedded in the model’s billions of parameters. Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand [6, 23, 24, 26, 40]. Indeed, there is a tension: an LLM may contain the facts needed for a task but still fail to recall them or apply them correctly when prompted. CoT prompting was one attempt to probe this internal knowledge by encouraging stepwise recall; however, CoT alone cannot guarantee retrieval of all relevant details. The model might skillfully reason with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge. This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial [47, 17]. To address these issues, researchers have proposed retrieval-augmented generation (RAG) techniques, which marry LLMs with external knowledge sources. Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence [44]. Lewis et al. [32] demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks. Borgeaud et al. [6] introduced the RETRO (Retrieval-Enhanced Transformer) model. A RETRO model of only 7 billion parameters matched the performance of a 178B parameter model without retrieval. He et al. [20] showed that an LLM with a retrieval step produces more faithful, grounded reasoning – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness. In general, augmenting LLMs with retrieval has been found to boost factual accuracy, reduce hallucinations, and expand the range of answerable queries. This is likely because LLMs gain new information during retrieval. LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data [49]. By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarity and difference.
@@
Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity [21, 13, 10, 11]. Contrastive methods work by pulling representations of semantically related inputs closer together while pushing those of unrelated inputs apart, fostering a latent space that reflects nuanced conceptual structure. While our evaluation does not involve contrastive training per se, it is motivated by a similar principle: if a model can consistently discern between distinct concepts, it demonstrates a deeper form of understanding—one likely to translate into stronger performance on downstream tasks.

### v3

@@ 5 Results
In this section we decompose the error rate into its components - how much of the error is due to missing information and how much is due to the inability to retrieve self-knowledge. Section 5.1 shows our main result. In the following section, we analyze how model size affects question quality and judgment accuracy (5.2).
@@ 5.1 LLMs Do Not Know What They Know
(a) LLM does not know what it has known: Self-talk helps improve judgment, but not as much as scientific QA. $p$-values: $p_{q_{1}}$ to $p_{q_{5}}$ = 0.000, $p_{q_{6}}$ = 0.013 (self vs. scientific answers).
(b) Scientific retrieval does not introduce more “content-level” new information for LLM to answer these questions.
Figure 4: Mechanisms of "LLMs do not know what they know".
Figure 4(a) presents our initial results from a random sample of 1000 patent pairs, examining how questions and answers affect classification performance. The baseline represents the initial accuracy of Llama 8B without any self-questioning or scientific retrieval. At this baseline level, Llama 8B incorrectly classifies approximately 70% of patent pairs as identical.
The blue line demonstrates the impact of adding questions alone. This modest improvement suggests that simply posing understanding questions helps the model perform better by providing structure that aids text parsing, though the effect of merely enumerating questions remains limited.
The orange line reveals that adding self-generated answers significantly improves performance. This substantial gain implies that the model already contains useful knowledge for the differentiation task, but this knowledge is not being effectively utilized. These results suggest that LLMs do not know what they know—bringing forth relevant information that enhances understanding can dramatically improve performance.
Finally, the green line shows the improvement achieved by providing retrieved answers. This additional performance gain indicates that while the model’s internal knowledge is valuable, it does not match the quality of complete information available in the training data. Notably, with six question-answer pairs, performance reaches the same level as the 405B parameter model, effectively eliminating the error rate.
Importantly, most of the error reduction is achieved through self-generated answers, suggesting that the model contains the most useful information within its internal knowledge representation. However, the model lacks the ability to effectively access and apply this knowledge when tackling tasks that require such domain expertise.
Our results demonstrate that models fail to retrieve the information necessary for the task, and the information they do retrieve is of lower quality than that obtained from additional sources. We next explore the factors driving this performance difference, considering two possible explanations. First, LLMs function as compressed representations of knowledge but do not preserve complete information. Scientific and technological information tends to be sparsely and unevenly distributed in training corpora, and the compression process significantly degrades the fidelity of such underrepresented knowledge. Second, LLMs may acquire new knowledge at inference time—meaning the external knowledge we provide is entirely unknown to the LLM and must be learned and incorporated during inference.
We argue that the first explanation is likely the primary cause of the improvement we observe. The models have likely already encountered the data we use: Llama 8B was released in July 2024444https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, when approximately 81% of the retrieved papers that generated scientific answers had already been published. However, while the models may have seen these papers during training, this exposure might not have significantly influenced the learned weights.
To further investigate what type of information is contained in inference-time retrieval versus the model’s internal knowledge, we analyze the answers by embedding all responses from both scientific sources and the LLM’s self-generated answers using the model from (Ghosh et al., 2024). To visualize and compare the distribution of embedding spaces, we reduce dimensionality to 2D and estimate boundaries using 5% Gaussian kernel density estimation—drawing contours that enclose the densest 95% of data points in each embedding set.
If LLMs substantially incorporate new scientific content during inference, we would expect their embedding regions to diverge meaningfully from those derived from self-generated responses. However, as shown in Figure 4 panel b, the embedding regions largely overlap. Notably, self-generated responses display a distinct outlier region, which we suspect reflects hallucinated content introduced by the LLM.
Despite occupying similar regions in the 2D embedding space that captures the greatest variance in the high-dimensional (768D) space, these two answer types differ in several meaningful ways. Scientific answers average $62.1\pm 1.8$ words and are presumably richer in detail and informativeness, drawing more heavily from scientific evidence. In contrast, self-generated answers are notably shorter, averaging only $38.2\pm 0.6$ words—approximately half the length of scientific responses. We therefore, interpert the results as suggesting that although the LLMs already have the knowledge needed to answer the question, and obtain better performances - this knowledge is compressed, less detailed and not as rich as the original scientific data.
5.2 It’s Not Only What to Ask, but Who Asks It
Self-questioning helps because it aligns the model’s attention with task-relevant concepts. This suggests a transferring question: are the most useful questions those that the same model would ask itself, or can questions authored by a different model—smaller or larger—better align the answerer’s knowledge with the task? We next test how question–answer sets travel across scale and where misalignment helps or hurts.
@@ 5.2 It’s Not Only What to Ask, but Who Asks It
The answer depends on which model is doing the answering. We test two models—LLaMA 3B and 8B—on questions generated by four others: LLaMA 1B, 8B, 70B, and 405B. We select 3B and 8B as answerer models because the 1B model produces low-confidence responses (94.4% with a confidence score of 0), indicating little understanding of content. In contrast, the 70B and 405B models already achieve strong performance without an external scientific context, leaving little room for improvement. The 3B and 8B models provide a more informative middle ground: they show some understanding but still benefit from superior prompts.
@@
As shown in Figure 5 panels a and b, we find that the 3B model performs best when answering questions from the 1B model, and the 8B model benefits from questions posed by both the 1B and 8B. This asymmetry reveals a critical insight: questions from larger models do not consistently help smaller models, and often introduce confusion. A more knowledgeable model cannot fully substitute for a smaller model’s understanding, suggesting that knowledge extraction is not smoothly transferable across model scales (Li et al., 2024). At the same time, we observe that questions from smaller models often align well with the answering capacity of larger models. This may be because larger models can readily interpret, build on, and give a better answer to smaller models’ questions. Large model questions appear to go beyond the bounds (and dimensionality) of smaller model understanding and answering capacity.
@@ 6 Final Remarks
Self-questioning not only enhances LLMs’ understanding but also reveals how they organize and access internal/external knowledge. While models hold useful internal insights, these are often latent and underused. Self-questioning helps activate this hidden compressed capacity, but falls short for complex tasks. Interestingly, the fundamental questions from smaller models often align with the needs of larger models. This suggests a scalable strategy: use smaller models to structure problems and guide deeper reasoning in stronger models.
Impact: our work and the associated new benchmark on patent distinction offer broad insights into how AI can be used to automate the evaluation of scientific and technical contributions.
@@ Appendix A Understanding Through Differentiation
Understanding is an ability that goes beyond merely recalling isolated facts or performing procedures; Grimm [2024] it fundamentally involves grasping the relationships that connect concepts and principles within a domain. This requires both an appreciation of and the ability to uncover underlying structures, such as causal links, correlations, hierarchies, or analogical parallels, which distinguish true comprehension from sophisticated pattern matching or associative recall. Evaluating an LLM’s understanding, therefore, requires methods that probe its capacity to navigate these relational structures, moving beyond tests of surface-level knowledge.
@@
Cognitive science, from [Gibson, 1969] through [Tversky, 1977] and Gentner’s structure-mapping work [Markman and Gentner, 2000] to recent relational-reasoning research [Alexander and Dumas, 2018], demonstrated that the ability to perceive, represent, and reason about crucial semantic relationships is built upon two complementary foundational processes: differentiation, contrasting, or “splitting”, and generalization, merging, or “lumping”. Differentiation is the capacity to distinguish concepts, entities, or situations, recognizing the features or relations that make them distinct – essentially, identifying what makes ’A’ different from ’B’. Without this ability to draw meaningful distinctions, even simple relationships cannot be properly conceptualized. Generalization, conversely, involves identifying commonalities across different instances, abstracting shared properties or underlying structures, and treating distinct items as equivalent for certain purposes, such as belonging to the same category or representing the same concept. These processes are fundamental not just for forming basic concepts, but for constructing the complex relational knowledge that constitutes deep understanding in any domain.
@@
Moreover, the ability to differentiate and generalize remains vital for constructing and refining our understanding of relationships between concepts, especially causal ones, which serve as the backbone of understanding. To explain why something occurs, we must first detect patterns: when does outcome $B$ follow event $A$, and when does it not? This requires the capacity to distinguish $A$ from alternatives like $A^{\prime}$, and to notice that $B$ follows $A$ but not $A^{\prime}$. Such contrasts form the basis for counterfactual reasoning, allowing us to imagine what would happen if everything about $A$ and $A^{\prime}$ were held constant except for one defining feature—if changing that feature flips the outcome, it signals a likely cause. At the same time, we must also recognize when observations are similar in respects that matter. Only by treating $A$ and $A^{\prime}$ as alike for relevant purposes, and then seeing whether their outcomes differ, can we meaningfully generalize and say that “$A$ causes BB.” This dual capacity—to detect when things are different and when they are effectively the same—is what underlies both the identification and the generalization of causal claims. Contemporary computational models and econometric frameworks reflect this structure explicitly, emphasizing the need for clear conceptual boundaries and controlled variation [Pearl, 2009, Lake et al., 2017, Heckman and Pinto, 2022]. Developmental research similarly shows that even young children rely on these kinds of structured comparisons to infer causes [Gopnik and Wellman, 2012].
@@
LLMs have demonstrated a remarkable ability to understand, interact with, and respond to human language Wang et al. [2018], Wang et al. [2019], OpenAI [2023], Brown et al. [2020]. Nevertheless, understanding language or recalling facts alone does not guarantee understanding higher-order concepts or the construction of a world model Vafa et al. [2024] that describes and explains how things relate to one another. The discussion above highlights that for LLMs to demonstrate comprehension of high-level concepts, they must be able to discern similarities and differences between complex concepts. Consequently, measuring an LLM’s proficiency in discerning similarity and difference within highly specific concepts and fields offers a powerful proxy for evaluating its understanding of the conceptual foundations and landscape of these fields.
@@
Our measure of understanding draws inspiration from previous benchmarks used to assess language comprehension in earlier generations of language models. GLUE Wang et al. [2018], for example, includes tasks that involve determining whether two sentences convey similar meanings or whether one entails the other. Similarly, our tasks focus on whether models can distinguish subtle conceptual differences. While prior benchmarks often relied on synthetic or crowd-sourced data, e.g. Bowman et al. [2015], our use of patent abstracts introduces a more challenging and high-stakes setting, requiring a fine-grained understanding of technical content.
@@ Appendix B Related Literature
Relations to other benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding. But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after [McIntosh et al., 2024, Wu et al., 2024, Davis, 2023]. High performance may reflect superficial pattern matching or memorization rather than true conceptual comprehension. In particular, current evaluation datasets frequently allow models to exploit statistical associations or training data overlaps to get the right answers without genuine understanding  [^5] https://www.sciencenews.org/article/ai-understanding-reasoning-skill-assess. Even when test content is novel, models can exploit formatting quirks. Alzahrani et al. [Alzahrani et al., 2024] showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions. Such brittleness indicates models are learning task-specific tricks (e.g., defaulting to a particular choice position or prompt wording) rather than demonstrating robust knowledge application. Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics [Arakelyan et al., 2024]. Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities [Burnell et al., 2023]. In sum, many current benchmarks are static and shortcut-prone, failing to truly assess whether an AI understands concepts or is simply leveraging surface regularities. Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase [Asthana et al., 2024] or contrast and counterfactual sets [Lewis and Mitchell, 2024]  [^6] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content [West et al., 2024]..
@@
Organizing internal and external knowledge: A core challenge in LLM understanding is how to activate the vast latent knowledge embedded in the model’s billions of parameters. Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand [Borgeaud et al., 2022, Hoang et al., , Hoang et al., 2023, Jaiswal et al., 2024, Pan et al., 2025]. Indeed, there is a tension: an LLM may contain the facts needed for a task but still fail to recall them or apply them correctly when prompted. The model might skillfully "reason" with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge. This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial [Schulhoff et al., 2024, Gopnik and Wellman, 2012]. To address these issues, researchers have proposed retrieval-augmented generation (RAG) techniques, which marry LLMs with external knowledge sources. Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence [Prince et al., 2024]. Lewis et al. [Lewis et al., 2020] demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks. Borgeaud et al. [Borgeaud et al., 2022] introduced the RETRO (Retrieval-Enhanced Transformer) model. A RETRO model of only 7 billion parameters matched the performance of a 178B parameter model without retrieval. He et al. [He et al., 2022] showed that an LLM with a retrieval step produces more faithful, grounded understanding – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness. In general, augmenting LLMs with retrieval has been found to boost factual accuracy, reduce hallucinations, and expand the range of answerable queries. This is likely because LLMs gain new information during retrieval. LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data [Shumailov et al., 2024]. By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarities and differences.
@@
Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity [He et al., 2020, Gao et al., 2021, Chen et al., 2020, Cohan et al., 2020]. Contrastive methods work by pulling representations of semantically related inputs closer together while pushing those of unrelated inputs apart, fostering a latent space that reflects nuanced conceptual structure. While our evaluation does not involve contrastive training per se, it is motivated by a similar principle: if a model can consistently discern between distinct concepts, it demonstrates a deeper form of understanding—one likely to translate into stronger performance on downstream tasks.